# CompdVision: Combining Near-Field 3D Visual and Tactile Sensing Using a Compact Compound-Eye Imaging System

### 12 Dec 2023 v1

@@ IV-B2 Blob Detection
Our sensor utilizes the HSV (Hue, Saturation, Value) color space adjustment technique to enhance the isolation of markers. The HSV color model is proficient in representing an image’s dominant wavelength of light, saturation, and brightness. By fine-tuning the HSV values, we can effectively mitigate specific components that could potentially impede accurate marker detection. Subsequently, the SimpleBlobDetector from OpenCV [27] is employed for the detection process within the masked regions, as shown in Fig. 5(c). This method allows for the accurate identification of markers and the extraction of their centroid coordinates.
@@ IV-B3 Marker Displacement Extraction
After the initial image capture, all markers are precisely detected for tracking. For each frame that follows, the detected blob points are compared and matched with those from the initial image, establishing correspondences based on nearest neighbors. These displacements represent contact deformation, thereby generating a 2D flow. The result of this processing is illustrated in Fig. 5(d). The 2D flow is transformed into a 6$\times$6 regular grid by interpolating the 2D displacements of the sparse points, a process facilitated by a function from the SciPy package [28]. This step enables the creation of a continuous field from markers, regardless of their sparse distribution or irregular spacing.
@@ V Experiments
In this section, we detail three experiments. The first experiment was dedicated to the training and evaluation of the stereo units in our sensor. The second experiment involved the training and evaluation of contact force measurement, which pertains to the tactile units of our sensor. These two experiments rigorously trained and evaluated the performance of our sensor’s fundamental units. In the third experiment, we tracked the evolution of 3D visual and tactile signals throughout a grasping task, providing an in-depth illustration of the operational combination between our sensor’s stereo and tactile units.
@@
Figure 7: Evaluation results of stereo depth estimation.
V-A Evaluation of Stereo Depth Estimation
We trained and evaluated the performance of our stereo vision units in depth estimation. We employed a series of performance metrics—fill rate, Z-accuracy, RMSE (Root Mean Square Error), and temporal noise—to evaluate the system’s efficacy [29]. A thorough evaluation of the performance of both the left and right stereo units was carried out.
We used a setup featuring a textured target and an electrical linear stage to generate a comprehensive dataset, as illustrated in Fig. 6. The textured target served as the object for depth mapping. The ground truth for depth was established by moving the electrical linear stage, which operated within a 0 to 70 mm range from the sensor surface. For each 1 mm interval within this range, we collected 20 images. The collected dataset was divided into two subsets: 60% for training and 40% for evaluation.
For the training process, we employed the dual annealing algorithm, using the training set to train the model. This global optimization method [30] was well-suited to our complex, multi-modal cost function, which involves multiple parameters for the SGBM algorithm and the WLS filter. The algorithm’s robustness in initializing conditions and its ability to escape local minima make it an excellent choice for high-dimensional optimization problems like ours.
@@ V-A1 Fill Rate
The fill rate metric quantifies the percentage of pixels in the depth map for which successful depth calculation was made. This was determined by defining an ROI that excludes the marker area and the boundary regions where disparity calculations are inherently limited due to the lack of corresponding pixels in the adjacent image. The fill rate was then calculated as the ratio of pixels with valid depth values to the total number of pixels within the ROI. It is noteworthy that the median fill rate for both the left and right stereo units attained a value of 100%.
@@ V-A2 Z-Accuracy
The Z-Accuracy metric assesses the accuracy of per-pixel depth estimation relative to the ground truth (GT). In this context, the GT is defined as the sum of the flange focal distance (distance from the image plane to the sensor surface) and the distance from the textured target to the sensor surface. The absolute difference between the GT and the collected depth map from the evaluation set was calculated. This computation was made relative to the best fitted plane to minimize the impact of potential errors in camera positioning, thereby generating a depth error map. Mathematically, the Z-Accuracy is expressed as the median value of these differences:
@@
The evaluation results demonstrated a consistent Z-Accuracy for both left and right stereo units, with higher accuracy observed at closer distances from the textured target.
@@ V-A3 RMSE
The RMSE metric assesses depth spatial noise in relation to the best fitted plane. Lower RMSE values indicate superior performance and heightened accuracy in depth estimation. The evaluation results demonstrated that the median RMSE for both left and right stereo units remained below 1.2% within a 70mm range from the sensor surface, signifying an effective depth estimation performance.
@@ V-A4 Temporal Noise
V-B Evaluation of Force Measurement
We used a learning-based method to train and evaluate the estimation of tangential and normal forces. This was achieved by analyzing the displacement field of the markers on the elastomer, a method that has proven to be robust [5][20].
Fig. 8 illustrates our experimental setup, which incorporates a 3-axis linear stage to control the sensor movement. We affixed a spherical 3D-printed indenter to the ATI Gamma Force and Torque (F/T) sensor. This setup allowed us to accurately record and measure the ground truth force as the sensor was pushed against the indenter with varying force amounts, directions, and locations on the sensor surface. The sensor’s interaction with the indenter was controlled meticulously. We pushed the sensor against the center and four surrounding points on the sensor surface. At each location, we moved down, left, and right with a consistent step size. This procedural variation introduced tangential and normal force, with the tangential force ranging from -1.16N to 1.27N, and the normal force ranging from 0N to 1.92N. We thus created a dataset containing 3500 measurements.
After data collection, we divided the dataset randomly, allocating 70% for training and the remaining 30% for testing. A Convolutional Neural Network (CNN) was utilized to train our data. The structured grid of the displacement field of the markers on the elastomer was inputted to the CNN model, which comprised two convolutional layers, three fully connected layers, and an output layer. Our CNN model was trained to predict the three forces $[F_{x},F_{y},F_{z}]$ where $F_{x}$ and $F_{y}$ represent the tangential forces, and $F_{z}$ represents the normal force. The results, as shown in Table II, indicate the RMSE for the force measurement on all three axes. The experiments and the results attest to our sensor’s robust capacity for contact force measurement.
@@ V-B Evaluation of Force Measurement
TABLE II: Evaluation Results of Force Measurement
@@
RMSE (N) 0.15 0.15 0.19
[…]
Figure 9: (a) The gripper approached the object. (b) The sensor made tactile contact with the objects. (c) The object was elevated. (d) The object was released from the grasp. (e) 3D visual and tactile sensing signals were captured throughout the process.
V-C Sensing in Grasping Tasks
The grasping task, depicted in Fig. 9(a-d), followed a sequential process. During the entire process, our algorithm ensured the simultaneous capture of 3D visual and tactile signals, as demonstrated in Fig. 9(e). The algorithm was capable of maintaining an online computation frequency of approximately 15 Hz, facilitated by a computer powered by an i9-10900 CPU.
The task began with the gripper’s approach towards the objects (Fig. 9(a)). This phase was dominated by visual signals, serving as the primary feedback for tracking the gripper’s progressive approach. Subsequently, the sensor established tactile contact with the objects (Fig. 9(b)). Here, visual signals receded, and tactile information became predominant, marked by the force exerted by the gripper. This was followed by the object’s elevation (Fig. 9(c)), where the gripper ensured a secure yet cautious hold to lift the object. Finally, the object was released (Fig. 9(d)). This phase was characterized by fading tactile signals and the re-emergence of visual signals, indicating the gripper’s expansion.
These signals offered rich insights into the interactions, recording the experiment comprehensively from start to finish.
VI Conclusion and Future Work
CompdVision represents a significant advancement in robotic sensing, providing a compact solution that combines near-field 3D visual and tactile modalities. The sensor offers the potential to tackle challenges in robotic applications, such as object recognition, navigation, and complex manipulation tasks. Its unique configuration allows for a seamless combination of 3D visual and tactile modalities, setting a new standard for multi-modal sensing in robotics.
Future research will focus on enhancing the sensor’s adaptability and performance in dynamic environments. We aim to use advanced methodologies and algorithms to improve the sensor’s responsiveness, precision, and durability. Additionally, the application of this sensor in robotic manipulation tasks will be explored. The objective is to understand how our sensor can improve the efficiency and autonomy of robotic manipulation, expanding its use beyond current applications.
Acknowledgement
The authors acknowledge the Nanosystem Fabrication Facility (NFF) of the Hong Kong University of Science and Technology and appreciate Mr. Mark Ellwood for proofreading and English editing.

### 15 Mar 2024 v2

@@ IV-B2 Blob Detection
Our sensor utilizes the HSV (Hue, Saturation, Value) color space adjustment method to enhance the isolation of markers. By fine-tuning the HSV filter values, we can mitigate specific components that could potentially impede marker detection. Subsequently, the SimpleBlobDetector from OpenCV [28] is employed for the detection process within the masked regions, as shown in Fig. 4(c). This method allows for the accurate identification of markers and the extraction of their centroid coordinates.
@@ IV-B3 Marker Displacement Extraction
After the initial image capture, all markers are precisely detected for tracking. For each frame that follows, the detected blob points are compared and matched with those from the initial image, establishing correspondences based on nearest neighbors. These displacements represent contact deformation, thereby generating a 2D flow. The result of this processing is illustrated in Fig. 4(d). The 2D flow is converted into a 6x6 grid through cubic interpolation of sparse point displacements, utilizing a SciPy function [29]. This step enables the creation of a continuous field from markers, regardless of their sparse distribution or irregular spacing.
IV-C Simultaneous 3D Visual and Tactile Sensing
The sensor can process both 3D visual and tactile data from a single image, owing to the compound-eye imaging system’s multiple vision units, each assigned to a specific task. The dual-focus microlens array within the system enables simultaneous 3D visual and tactile sensing without requiring modality conversion. By using multi-threading on a computer with an i9-10900 CPU, the sensor achieves a sensing rate of 25 frames per second. This method significantly enhances operational efficiency and responsiveness.
@@ V Experiments
In this section, we detail two experiments to calibrate and evaluate the performance of depth estimation and force measurement.
V-A Depth Estimation Calibration
We trained and evaluated the performance of our stereo depth estimation. We employed a series of performance metrics—fill rate, Z-accuracy, Root Mean Square Error (RMSE), and temporal noise—to evaluate the system’s efficacy [30]. A thorough evaluation of the performance of both the left and right stereo pairs was carried out.
We used a setup featuring a textured target and an electrical linear stage to generate a dataset, as illustrated in Fig. 5. The textured target served as the object for depth mapping. The ground truth for depth was established by moving the electrical linear stage, which operated within a 0 to 70 mm range from the sensor surface. For each 1 mm interval within this range, we collected 20 images. The collected dataset was divided into two subsets: 60% for training and 40% for evaluation.
@@ V-A Depth Estimation Calibration
Figure 6: Evaluation results of depth estimation.
For the training process, we employed the dual annealing algorithm, using the training set to train the model. This global optimization method [31] was well-suited to our complex, multi-modal cost function, which involves multiple parameters for the SGBM algorithm and the WLS filter.
@@ V-A1 Fill Rate
The fill rate metric measures the success rate of depth calculation as a percentage of pixels in the depth map, excluding marker areas and boundary regions where disparity calculation is limited. Calculated as the ratio of pixels with valid depth to total ROI pixels, both left and right stereo pairs achieved a 100% mean fill rate.
@@ V-A2 Z-Accuracy
The Z-Accuracy metric evaluates the accuracy of depth estimation per pixel compared to the ground truth (GT), defined as the sum of the flange focal distance and the distance from the target to the sensor. We calculated the absolute difference between GT and the measured depth, relative to the best-fitted plane, to reduce errors from camera positioning, resulting in a depth error map. Mathematically, the Z-Accuracy is expressed as the median value of these differences:
@@
The evaluation results demonstrated consistent Z-Accuracy for both left and right stereo pairs, with higher accuracy observed at closer distances to the textured target. The standard deviations (STDs) are below 0.23mm.
@@ V-A3 RMSE
The metric evaluates spatial noise in depth measurements relative to the best-fitted plane. Lower RMSE values denote superior performance and increased accuracy in depth estimation.
Results show that mean RMSE and STD values remain below 1.2% and 0.51%, respectively. However, these values tend to rise when the target nearly obscure the sensor surface, primarily due to reduced light conditions and the lack of internal lighting. Despite this, the sensor’s performance remains effective. This is largely due to the white sensor shell, which allows external light to penetrate, and the microlens array’s low light intensity requirements [21]. Consequently, reliable depth estimation is ensured even for objects in close proximity to the sensor surface without internal lighting.
@@ V-A4 Temporal Noise
V-B Force Measurement Calibration
We used a learning-based method to train and evaluate the estimation of tangential and normal forces. This was achieved by analyzing the displacement field of the markers, a method that has proven to be robust [5][12].
Fig. 7 illustrates our experimental setup, incorporating a 3-axis linear stage to control sensor movement. We affixed 3D-printed indenters of five different shapes (triangle, rectangle, hexagon, circular column and sphere) to the ATI Gamma Force and Torque (F/T) sensor. This setup allowed us to record and measure the ground truth force as the sensor was pressed against the various indenters with varying force amounts, directions, and locations on the sensor surface. The sensor’s interaction with indenters was controlled automatically. Indentors were pressed against the center and four surrounding points on sensor surface. For each indentation, a step size of 0.025 mm was used to incrementally increase the indentation depth up to a maximum of 4mm to introduce normal force. At each indentation step of 0.5mm, to introduce tangential force, additional movements were made up, down, left, and right by up to 3mm, each movement executed with a step size of 0.2 mm. This procedure resulted in the introduction of both tangential and normal forces, with tangential force values ranging from -1.83N to 2.04N, and normal force values ranging from 0N to 4.85N. Consequently, a dataset comprising 16,000 measurements was created.
After collecting data, we randomly split the dataset, allocating 70% for training and 30% for testing. Our data processing employed a Convolutional Neural Network (CNN) model with multiple convolutional and fully connected layers. The model incorporates five convolutional layers and batch normalization for stable learning. Enhanced by max pooling and dropout for regularization, the architecture concludes with fully connected layers, leading to three output neurons for the forces $[F_{x},F_{y},F_{z}]$, where $F_{x}$ and $F_{y}$ denote tangential forces, and $F_{z}$ denotes the normal force. The evaluation results using the testing dataset, as shown in Table III, indicate the RMSE for the force measurement on all three axes. The experiments and results underscore our sensor’s robust capability for accurate contact force measurement.
@@ V-B Force Measurement Calibration
TABLE III: Evaluation Results of Force Measurement
@@
RMSE (N) 0.17 0.18 0.26
VI Discussion and Conclusion
CompdVision represents a significant advancement in robotic sensing, offering a compact solution for a multi-modal sensor capable of simultaneous 3D visual and tactile sensing. Its innovative design and configuration set a new standard for multi-modal sensing in robotics.
While the sensor demonstrates promising capabilities, a thorough understanding of its limitations is crucial. It demonstrates great depth estimation capabilities, marked by excellent Z-accuracy and RMSE, even under conditions where objects are in close proximity or obscure the sensor. However, its performance heavily depends on external lighting, a limitation that could be mitigated by incorporating internal lighting to improve low-light performance. Additionally, the sensor’s reliance on black markers may pose challenges in blob detection when interacting with black objects. By applying a special coating to the markers and customizing tactile units to detect markers against diverse backgrounds, we can enhance the sensor’s tactile sensing adaptability. Future research will focus on enhancing the sensor’s robustness and performance in dynamic environments and exploring its applications in robotic manipulation tasks.